### v1

Abstract
The interplay between topology and quantum criticality gives rise to the notion of symmetry-enriched criticality, which has attracted considerable attention in recent years. However, its non-Hermitian counterpart remains largely unexplored. In this Letter, we show how parity-time ($\mathcal{PT}$) symmetry enriches non-Hermitian critical points, giving rise to a topologically distinct non-unitary universality class. By analytically investigating non-Hermitian free fermion models with $\mathcal{PT}$ symmetry, we uncover a new class of conformally invariant non-unitary critical points that host robust topological edge modes. Remarkably, the associated topological degeneracy is surprisingly encoded in the purely imaginary part of the entanglement entropy scaling—a feature absent in Hermitian systems. The underlying mechanism for the emergence of edge states at non-Hermitian criticality is traced to a generalized mass inversion that is absent in Hermitian systems.
Introduction.—Non-Hermitian (nH) quantum systems have recently attracted considerable attention due to their unique properties beyond Hermitian counterparts [1, 2, 3, 4, 5] and are closely related to diverse experimental platforms [6, 7, 8, 9, 10]. The characterization of quantum phases in nH systems is now well established [11, 12, 13, 14, 15, 16], revealing distinctive phenomena such as the non-Hermitian skin effect [17, 18, 19, 20, 21, 22] and non-Hermitian topological phases [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. In contrast, phase transitions in nH systems—typically described by non-unitary conformal field theories (CFTs)—remain less understood, although they have recently drawn growing interest from the perspective of quantum entanglement [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70].
On a different front, the universality class of quantum phase transitions can be further enriched by global symmetries, giving rise to topologically distinct universality classes, now referred to as symmetry-enriched quantum criticality (SEQC) [71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88] or, more generally, gapless symmetry-protected topological (gSPT) states  [89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121]. The discovery of topological physics in quantum critical systems opens new avenues for classifying phase transitions within the same universality class, fundamentally enriching the textbook understanding of phase transitions and going beyond the traditional paradigms of statistical and condensed matter physics. Remarkably, the coexistence of symmetry-protected edge states with a gapless bulk gives rise to intriguing topological phenomena absent in gapped topological phases, including nontrivial conformal boundary conditions [76, 92], algebraically localized edge modes [71, 117], universal bulk-boundary correspondence [76, 102, 83], and intrinsically gapless topological phases [96, 115].
A fundamental open question is whether the notion of SEQC can be generalized to non-unitary criticality in non-Hermitian systems, and if so, what mechanisms govern such phenomena. Despite its importance, this issue has remained largely unexplored due to the intrinsic difficulty of systematically investigating quantum criticality in generic non-Hermitian many-body systems.
To make progress in answering the above questions, we systematically explore $\mathcal{PT}$-symmetry–enriched non-unitary criticality in a family of one-dimensional nH free-fermion models, uncovering a new class of non-unitary critical points that host robust topological edge modes. Using the combination of exact solution and numerical simulations, we show that although both critical points in these models are described by non-unitary CFTs with central charge c=2c=-2, one of them is further enriched by $\mathcal{PT}$ symmetry, giving rise to topologically non-unitary criticality with stable edge states. Remarkably, the associated topological degeneracies are encoded in the purely imaginary part of the entanglement entropy scaling, a feature unique to non-Hermitian systems. Finally, we elucidate the emergence of topological edge states at this critical point by proposing an entirely new mechanism—generalized mass inversion—that is absent in Hermitian counterparts.
Non-Hermitian SSH model and its $\alpha$-extensions—We begin with the nH SSH chain with balanced gain/loss. In the momentum space, the single-particle Hamiltonian is
$\mathcal{H}_{k}=\begin{pmatrix}iu&v_{k}\\ v_{k}^{*}&-iu\end{pmatrix},\qquad v_{k}=v-we^{-ik},$ (1)
where $k$ is the single-particle momentum, $v,w\in\mathbb{R}$ are the intra-/inter-cell hoppings, and $u\in\mathbb{R}$ is the staggered imaginary on-site potential (see Fig. 1(b) for a schematic). With parity $\mathcal{P}=\sigma_{x}$ and time reversal symmetry $\mathcal{T}=\mathcal{K}$, the model satisfies $\sigma_{x}\mathcal{H}_{k}^{*}\sigma_{x}=\mathcal{H}_{k}$, i.e., it is $\mathcal{PT}$-symmetric. The bulk dispersion $E_{k}=\pm\sqrt{|v_{k}|^{2}-u^{2}}$ is real when $\min_{k}|v_{k}|>u$ ($\mathcal{PT}$-symmetric) and becomes complex otherwise, signaling spontaneous $\mathcal{PT}$ breaking.
From the bulk-boundary correspondence, the number of edge-mode pairs under the open boundary condition (OBC) is directly related to the winding number computed in the momentum space [28]
$\omega=\frac{1}{2\pi}\int_{-\pi}^{\pi}\!dk\,\arg v_{k},$ (2)
which depends only on $v_{k}$. The vertical dashed line $w=v$ in Fig. 1(a) therefore marks a change in edge-mode count that is independent of $u$. Note that inside the $\mathcal{PT}$-broken wedge the dashed line is not a phase boundary, because the Berry (Zak) phase, which determines the phase structure, no longer corresponds to the winding number there [122].
Turning on $u$ has a genuinely nH effect on criticality. At $u=0$ the Hermitian SSH chain has a single quantum critical point (QCP) at $w=v$ (black dot). For $u>0$, this point splits into two straight critical lines $u=\pm(w-v)$ that border the $\mathcal{PT}$-broken wedge. The left critical line lies in the $\omega=0$ sector and is topologically trivial. The right critical line lies in $\omega=1$; it is topological and is the main focus of this work. We then extend the construction to models with richer topological content in this work.
To realize a broader class of nH QCPs with nontrivial topology, we generalize the nH SSH model using an $\alpha$-chain construction inspired by Hermitian counterparts [72]. Specifically, we define the $\alpha$-nH SSH model as a $\mathcal{PT}$-symmetric free-fermion chain with long-range hopping,
$v_{k}^{(\alpha)}=v\,e^{-i(\alpha-1)k}-w\,e^{-i\alpha k}.$ (3)
The case $\alpha=1$ recovers the nH SSH model. For $\alpha\geq 2$, long-range hopping preserves the three-region structure of Fig. 1(a) while increasing the edge-mode multiplicity. In the $\mathcal{PT}$-symmetric gapped regime |vw|>u|v-w|>u under OBC, the side $v>w$ hosts $\alpha-1$ edge-mode pairs (winding number $\omega=\alpha-1$), whereas the side $w>v$ hosts $\alpha$ pairs (winding number $\omega=\alpha$). Accordingly, there are two nH critical lines that lie in distinct winding sectors: the one with lower winding number ($\omega=\alpha-1$) and the one with higher winding number ($\omega=\alpha$). In the main text, we focus on the $\alpha=1$ case and include the $\alpha=2$ model as a representative higher $\alpha$ example (see Fig. 1(c) for a schematic).
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Figure 1: (a) Phase diagram of the $\mathcal{PT}$-symmetric nH SSH models in the (u,wv)(u,w-v) plane. The Hermitian critical point at $(0,0)$ splits into two non-Hermitian critical lines $u=\pm(w-v)$ that border the $\mathcal{PT}$-broken region. The dashed line $w=v$ separates winding sectors $\omega=0$ and $\omega=1$. (b) Schematic nH SSH chain: red/blue circles denote $\pm iu$; black/purple bonds denote Hermitian hoppings. (c) Schematic $\alpha=2$ generalization.
Detecting non-unitary CFT at non-Hermitian criticality.—We assess conformal criticality in the $\alpha$-nH SSH models via entanglement scaling and a complementary Casimir check. Using the correlation matrix method [122], the entropy for a subsystem of length $\ell_{A}$ is
$S_{A}=\sum_{n}\left[-\nu_{n}\log\nu_{n}-(1-\nu_{n})\log(1-\nu_{n})\right],$ (4)
[…]
First we review the ordinary ($\alpha=1$) nH SSH model. At its trivial QCP, the correlation eigenvalues are real and occur in $(\nu,1{-}\nu)$ pairs, ensuring that the entanglement entropy remains real and scales with central charge c=2c=-2, consistent with a non-unitary CFT (see Fig. 2(a)).
In contrast, the topologically nontrivial QCP exhibits a complex conjugate pair of correlation eigenvalues, $\nu_{\pm}=1/2\pm i\mathcal{I}$, which lead to an imaginary contribution to the entanglement entropy. While previous studies [36] encountered ambiguities due to the multivalued complex logarithm, we restore the correct analytic structure by consistently shifting one logarithmic branch by $2\pi$. This yields an additional term $i\pi$ in the entropy, while preserving the scaling form with c=2c=-2 in the real part (see Fig. 2(b)). Unlike in Hermitian systems, the entropy here acquires a constant imaginary part; we defer its physical interpretation to a later section. Full details of the branch-cut prescription are provided in the Supplementary Material [122].
We now turn to the $\alpha=2$ model to examine how these features generalize. At the QCP with higher winding number ($\omega=2$), the entanglement spectrum contains two complex-conjugate eigenvalue pairs. Applying the same branch-cut prescription, the entropy acquires an additional term $2i\pi$, while maintaining the c=2c=-2 scaling (see Fig. 2(c)). The doubled imaginary part reflects the contribution from both complex pairs.
A complementary signature of criticality comes from the finite-size scaling of the ground-state energy under periodic boundary condition (PBC). In CFT, the ground-state energy takes the form
$E_{0}^{\mathrm{PBC}}(L)=L\epsilon-\frac{\pi v_{F}c}{6L}+\cdots,$ (5)
where $\epsilon$ is the bulk energy density and $v_{F}$ is the Fermi velocity.
For all $\alpha$-nH SSH models, the bulk dispersion remains identical across phases, enabling a unified scaling analysis near the critical points. As shown in Fig. 2(d), the finite-size corrections to the ground-state energy exhibit clear $1/L$ behavior with a coefficient consistent with c=2c=-2, corroborating the entanglement entropy results. Additional results for OBC and higher-$\alpha$ models—covering both entanglement entropy and ground-state energy—are fully consistent with c=2c=-2 and presented in the SM. Together, these findings demonstrate that all critical points in the $\alpha$-nH SSH models are described by a non-unitary CFT with central charge c=2c=-2.
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
Figure 2: Entanglement entropy $S_{A}$ as a function of the subsystem size $\ell_{A}$ for a total system length $L=10000$. Blue dots indicate the real part of $S_{A}$, red dots indicate the imaginary part, and the blue line shows the numerical fit. (a) $\alpha=1$, $(v,w,u)=(2,1,1-10^{-12})$ — trivial QCP, $\operatorname{Re}[S_{A}]=-0.665\ln[\sin(\pi\ell_{A}/L)]-10.18$. (b) $\alpha=1$, $(v,w,u)=(1,2,1-10^{-12})$ — topological QCP, $\operatorname{Re}[S_{A}]=-0.665\ln[\sin(\pi\ell_{A}/L)]-10.87$. (c) $\alpha=2$, $(v,w,u)=(1,2,1-10^{-12})$ — higher winding number QCP, $\operatorname{Re}[S_{A}]=-0.665\ln[\sin(\pi\ell_{A}/L)]-11.35$. (d) Finite-size scaling of the ground-state energy in the $\alpha$-nH SSH model under PBC with $v_{F}=\sqrt{2}$. All results support that these QCPs are described by a non-unitary CFT with central charge c=2c=-2.
Entanglement–edge correspondence and the imaginary entropy—At the critical points of the $\alpha$-nH SSH chain, we observe a correspondence between the physical and entanglement spectra. Under OBC, the system supports localized edge modes with purely imaginary energies $E=\pm iu$. Under PBC, the entanglement spectrum features complex-conjugate eigenvalue pairs of the form $\nu=1/2\pm i\mathcal{I}$, whose number matches that of the OBC edge modes [see Fig. 3(a,b)]. Using the standard correlation–entanglement mapping $\nu=(1+e^{\epsilon})^{-1}$, equivalently
$\epsilon=\log\frac{1-\nu}{\nu}=\pm 2i\arctan(2\mathcal{I}),$ (6)
corresponding to a purely imaginary entanglement energy. This generalizes the Li–Haldane picture: although both energy are no longer real or degenerate, the one-to-one counting and the vanishing real parts ($\operatorname{Re}[E]=\operatorname{Re}[\epsilon]=0$) still match.
Refer to caption(a)
Refer to caption(b)
Figure 3: Entanglement spectrum of the $\alpha$–nH SSH model in the topological phase. The left panels show the real part $\operatorname{Re}[\nu]$ and the right panels show the imaginary part $\operatorname{Im}[\nu]$ of the eigenvalues. (a) For $\alpha=1$, a single complex-conjugate pair is present. (b) For $\alpha=2$, two complex-conjugate pairs appear. In both cases, the number of complex-conjugate pairs matches the number of edge modes.
A key consequence concerns the imaginary part of the entanglement entropy. Requiring the correct c=2c=-2 scaling at all critical points forces us to adopt the same branch-cut convention used in the previous section. With this convention, each complex-conjugate pair contributes $i\pi$ to $S_{A}$. Together with the entanglement spectrum–edge correspondence, at a QCP with winding number $\omega$,
$S_{A}=-\frac{2}{3}\log\ell_{A}-i\pi\omega+\mathrm{const.}$ (7)
This identifies a subleading imaginary contribution $i\pi\omega$. Unlike in Hermitian SEQC, where topological features at criticality are reflected only in the entanglement spectrum [101, 83], the nH case renders them directly visible in the imaginary part of the entanglement entropy—a feature unique to nH systems.
It is then natural to extend the same branch choice into the adjacent SPT phase by continuity. With that extension, $\operatorname{Im}[S_{A}]$ takes the same constant value throughout each fixed-$\omega$ sector. Thus $\operatorname{Im}S_{A}$ serves as a topological index reflecting the winding number $\omega$ across the $\mathcal{PT}$-symmetric sector.
$\mathcal{PT}$ symmetry-protected topological edge states at non-unitary criticality.—The $\alpha$-nH SSH chain supports gapless topological features at criticality. We now show that these features are protected by $\mathcal{PT}$ symmetry, and that the protection is lost once $\mathcal{PT}$ symmetry is broken. From the dispersion $E_{k}=\pm\sqrt{|v_{k}|^{2}-u^{2}}$ the $\mathcal{PT}$-symmetric regime is characterized by the requirement $|v_{k}|\geq u$ for all $k$. In particular, this enforces $v_{k}\neq 0$ throughout the Brillouin zone, so the loop traced by $v_{k}$ never crosses the origin and the winding number $\omega$ is pinned. Equivalently, the number of OBC edge-mode pairs is fixed within a $\mathcal{PT}$-symmetric sector; this description includes the QCPs themselves.
[…]
To further test the stability of the edge structure, we add moderate disorder to both the hopping amplitudes and the on-site potentials while preserving global $\mathcal{PT}$ symmetry and keeping the system at criticality. As shown in Fig. 4(a), the imaginary part of the entanglement entropy, $\operatorname{Im}[S_{A}]$, remains the same across all disorder realizations. This indicates that the topological structure is robust against disorder.
Together, these results demonstrate that the generalized $\alpha$-nH SSH model hosts a robust topological edge states protected by $\mathcal{PT}$ symmetry, independent of a bulk gap or translational invariance. The critical point thus realizes a $\mathcal{PT}$ symmetry-enriched non-unitary CFT.
Refer to caption(a)
Refer to caption(b)
Figure 4: (a) Disorder-averaged entanglement entropy $S_{A}$ at its topological QCP ($v(x)=1+\delta(x)$, $w=2$, $u(x)=1-\delta(x)-10^{-10}$; $L=1000$; $\delta(x)\in[-0.999,0.999]$). Averaged over 10001000 disorder realizations; error bars denote ±1 s.e.m. across realizations. Blue dots represent the real part of $S_{A}$, red dots represent the imaginary part. (b) Interface schematic between a trivial gapped region (left) and the nH SSH chain at its topological QCP (right). Shown are the mass profile $m(x)$ (black), the spectral gap $\Delta(x)=\sqrt{m(x)^{2}-u(x)^{2}}$ (blue), and the bound-state density $|\psi(x)|^{2}$ (red) localized at the interface.
Generalized mass inversion in $\mathcal{PT}$ symmetry-enriched non-Hermitian QCP—A distinctive feature of the $\alpha$-nH SSH model is the appearance of an intermediate $\mathcal{PT}$-broken symmetry-breaking phase absent in Hermitian free-fermion systems. This additional phase shifts the location of the critical point and is closely tied to the persistence of edge modes across the transition. In the Hermitian $\alpha\geq 2$-chain, topological edge states rely on kinetic inversion with long-range hopping to maintain edge states at criticality. By contrast, in the nH SSH model, the survival of edge modes follows from a simple generalization of mass inversion without requiring long-range hopping ($\alpha=1$).
Near the critical point, the low-energy continuum Hamiltonian takes the Dirac-like form
$\mathcal{H}=\begin{bmatrix}iu&-\partial_{x}+m\\ \partial_{x}+m&-iu\end{bmatrix}$, (8)
where $m$ is the mass profile and $u$ the non-Hermitian strength. The eigenvalue equations
$(-\partial_{x}+m)\psi_{B}(x)$ $=(E-iu)\psi_{A}(x),$ (9)
$(\partial_{x}+m)\psi_{A}(x)$ $=(E+iu)\psi_{B}(x)$ (10)
decouple at $E=\pm iu$, yielding localized edge solutions with decay length $\sim 1/m$ for $m>0$ even as the bulk gap $\sqrt{m^{2}-u^{2}}$ closes.
More generally, this is due to when we have more than one mass operators $\{m_{a}\Gamma_{a}\}$ with $\Gamma_{a}$ being the gamma matrices in the Dirac-like Hamiltonian. Here some of the $m_{a}$ can be imaginary. Edge localization is set by the domain-wall mass—the component that flips across the boundary—so $\xi\sim 1/|m_{bdy}|$. The bulk gap, however, is controlled by the spectral mass $\Delta=\sqrt{\sum_{a}m_{a}^{2}}$. Thus these scales are able to decouple when we have multiple mass term. Including this scale decoupling together with ordinary mass inversion is what we refer to as generalized mass inversion.
This mechanism is absent in Hermitian systems, where the gap is always enhanced and cannot close without delocalizing the edge modes. This makes a free-fermion SEQC without long-range hopping possible only in the nH setting. For higher $\alpha$, a generalized mass inversion coexists with kinetic inversion, see the SM [122].
The schematic in Fig. 4(b) depicts an interface geometry (left: trivial gapped phase; right: nH SSH at its topological QCP). Unlike the OBC case, the interface equations remain coupled and lack a simple closed-form solution; nevertheless, generalized mass inversion pins an edge state at the interface [122].
Discussion and concluding remarks.—To summarize, we uncover a new class of non-unitary criticality enriched by $\mathcal{PT}$ symmetry, which hosts robust topological edge states. Specifically, we construct a broad family of one-dimensional $\mathcal{PT}$-symmetric nH free-fermion models with $\alpha$-range hopping, all exhibiting critical points described by non-unitary CFTs with central charge c=2c=-2, as confirmed numerically through the scaling of entanglement entropy and ground-state energy. More importantly, we unambiguously demonstrate that these nH critical points support nontrivial edge states protected by $\mathcal{PT}$ symmetry. The associated topological degeneracy is reflected in the bulk entanglement spectrum—representing a generalized Li-Haldane correspondence at nH criticality—and, remarkably, is also encoded in the imaginary part of the entanglement entropy, a feature unique to nH systems. Finally, we analytically show that the topological edge modes at these QCPs are stabilized by a new mechanism, generalized mass inversion—without requiring long-range hopping—highlighting a distinct protection mechanism from Hermitian counterparts.
Looking ahead, identifying symmetry-enriched nH criticalities in higher dimensions and in interacting many-body systems represents an interesting direction for future work. Regarding experimental realization, we note that nH free-fermion lattice models and entanglement-based observables can be implemented and probed using quantum and classical simulators, including ultracold atoms [123, 124], superconducting quantum processors [125, 126], and phononic systems [127]. Our work not only generalizes the notion of topology to the realm of non-unitary criticality but also contributes to a deeper understanding of critical phenomena in nH systems.
Note added: After completing this manuscript, we became aware of a related independent study on the critical edge states in one-dimensional nH free-fermion chains [128].
Acknowledgement: We thank Xueda Wen for helpful discussion. K.-H. Chou and P.-Y. Chang was supported by National Science and Technology Council of Taiwan under Grants No. NSTC 113-2112-M- 007-019, 114-2918-I-007-015. X.-J. Yu was supported by the National Natural Science Foundation of China (Grant No.12405034) and a start-up grant from Fuzhou University. P.-Y.C acknowledges support from RIKEN Center for Interdisciplinary Theoretical and Mathematical Sciences and National Center for Theoretical Sciences, Physics Division. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science, and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities.

### v2

@@
 By analytically investigating non-Hermitian free fermion models with $\mathcal{PT}$ symmetry, we uncover a new class of conformally invariant non-unitary critical points that host robust topological edge modes.
-Remarkably, the associated topological degeneracy is surprisingly encoded in the purely imaginary part of the entanglement entropy scaling—a feature absent in Hermitian systems.
+Remarkably, the associated topological degeneracies are surprisingly encoded in the purely imaginary part of the entanglement entropy scaling—a feature absent in Hermitian systems.
 The underlying mechanism for the emergence of edge states at non-Hermitian criticality is traced to a generalized mass inversion that is absent in Hermitian systems.
@@
 The characterization of quantum phases in nH systems is now well established [11, 12, 13, 14, 15, 16], revealing distinctive phenomena such as the non-Hermitian skin effect [17, 18, 19, 20, 21, 22] and non-Hermitian topological phases [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
-In contrast, phase transitions in nH systems—typically described by non-unitary conformal field theories (CFTs)—remain less understood, although they have recently drawn growing interest from the perspective of quantum entanglement [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70].
-On a different front, the universality class of quantum phase transitions can be further enriched by global symmetries, giving rise to topologically distinct universality classes, now referred to as symmetry-enriched quantum criticality (SEQC) [71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88] or, more generally, gapless symmetry-protected topological (gSPT) states  [89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121].
+In contrast, phase transitions in nH systems—typically described by non-unitary conformal field theories (CFTs)—remain less understood, although they have recently drawn growing interest from the perspective of quantum entanglement [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74].
+On a different front, the universality class of quantum phase transitions can be further enriched by global symmetries, giving rise to topologically distinct universality classes, now referred to as symmetry-enriched quantum criticality (SEQC) [75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92] or, more generally, gapless symmetry-protected topological states  [93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125].
 The discovery of topological physics in quantum critical systems opens new avenues for classifying phase transitions within the same universality class, fundamentally enriching the textbook understanding of phase transitions and going beyond the traditional paradigms of statistical and condensed matter physics.
-Remarkably, the coexistence of symmetry-protected edge states with a gapless bulk gives rise to intriguing topological phenomena absent in gapped topological phases, including nontrivial conformal boundary conditions [76, 92], algebraically localized edge modes [71, 117], universal bulk-boundary correspondence [76, 102, 83], and intrinsically gapless topological phases [96, 115].
-A fundamental open question is whether the notion of SEQC can be generalized to non-unitary criticality in non-Hermitian systems, and if so, what mechanisms govern such phenomena.
-Despite its importance, this issue has remained largely unexplored due to the intrinsic difficulty of systematically investigating quantum criticality in generic non-Hermitian many-body systems.
-To make progress in answering the above questions, we systematically explore $\mathcal{PT}$-symmetry–enriched non-unitary criticality in a family of one-dimensional nH free-fermion models, uncovering a new class of non-unitary critical points that host robust topological edge modes.
-Using the combination of exact solution and numerical simulations, we show that although both critical points in these models are described by non-unitary CFTs with central charge c=2c=-2, one of them is further enriched by $\mathcal{PT}$ symmetry, giving rise to topologically non-unitary criticality with stable edge states.
-Remarkably, the associated topological degeneracies are encoded in the purely imaginary part of the entanglement entropy scaling, a feature unique to non-Hermitian systems.
+Remarkably, the coexistence of symmetry-protected edge states with a gapless bulk gives rise to intriguing topological phenomena absent in gapped topological phases, including nontrivial conformal boundary conditions [80, 96], algebraically localized edge modes [75, 121], universal bulk-boundary correspondence [80, 106, 87, 126], and intrinsically gapless topological phases [100, 119].
+A crucial open question is how the concept of SEQC manifests in nH systems.
+Progress on this front remains challenging because studying generic nH many-body systems is notoriously difficult, plagued by numerical instabilities and inaccuracies that obscure the underlying physics [127, 128, 16, 68, 129, 130].
+To make substantive progress, we adopt a strategy centered on exactly solvable models.
+We systematically explore $\mathcal{PT}$-symmetry–enriched non-unitary criticality in a family of one-dimensional nH free-fermion models, uncovering a new class of non-unitary critical points that host robust topological edge modes.
+Using the combination of exact solution and numerical simulations, we show that although all critical points in these models are described by non-unitary CFTs with central charge c=2c=-2, at least one of them is further enriched by $\mathcal{PT}$ symmetry, giving rise to topologically non-unitary criticality with stable edge states.
+Remarkably, the associated topological degeneracies are encoded in the purely imaginary part of the entanglement entropy scaling, a feature unique to nH systems.
 Finally, we elucidate the emergence of topological edge states at this critical point by proposing an entirely new mechanism—generalized mass inversion—that is absent in Hermitian counterparts.
@@
 1(a) therefore marks a change in edge-mode count that is independent of $u$.
-Note that inside the $\mathcal{PT}$-broken wedge the dashed line is not a phase boundary, because the Berry (Zak) phase, which determines the phase structure, no longer corresponds to the winding number there [122].
+Note that inside the $\mathcal{PT}$-broken wedge the dashed line is not a phase boundary, because the Berry (Zak) phase, which determines the phase structure, no longer corresponds to the winding number there [132, 28, 133].
 Turning on $u$ has a genuinely nH effect on criticality.
@@
 First we review the ordinary ($\alpha=1$) nH SSH model.
-At its trivial QCP, the correlation eigenvalues are real and occur in $(\nu,1{-}\nu)$ pairs, ensuring that the entanglement entropy remains real and scales with central charge c=2c=-2, consistent with a non-unitary CFT (see Fig.
+At QCP wv=uw-v=-u, the correlation eigenvalues are real and occur in $(\nu,1{-}\nu)$ pairs, ensuring that the entanglement entropy remains real and scales with central charge c=2c=-2, consistent with a non-unitary CFT (see Fig.
 2(a)).
-In contrast, the topologically nontrivial QCP exhibits a complex conjugate pair of correlation eigenvalues, $\nu_{\pm}=1/2\pm i\mathcal{I}$, which lead to an imaginary contribution to the entanglement entropy.
+In contrast, the QCP wv=uw-v=u exhibits a complex conjugate pair of correlation eigenvalues, $\nu_{\pm}=1/2\pm i\mathcal{I}$, which lead to an imaginary contribution to the entanglement entropy.
 While previous studies [36] encountered ambiguities due to the multivalued complex logarithm, we restore the correct analytic structure by consistently shifting one logarithmic branch by $2\pi$.
@@
 Unlike in Hermitian systems, the entropy here acquires a constant imaginary part; we defer its physical interpretation to a later section.
-Full details of the branch-cut prescription are provided in the Supplementary Material [122].
+Full details of the branch-cut prescription are provided in the Supplementary Material (SM) [133].
 We now turn to the $\alpha=2$ model to examine how these features generalize.
@@
 All results support that these QCPs are described by a non-unitary CFT with central charge c=2c=-2.
-Entanglement–edge correspondence and the imaginary entropy—At the critical points of the $\alpha$-nH SSH chain, we observe a correspondence between the physical and entanglement spectra.
+Entanglement–edge correspondence and the imaginary entropy—We now focus on the topological aspects of these non-unitary critical points.
+Specifically, at the critical points of the $\alpha$-nH SSH chain, we observe a correspondence between the physical and entanglement spectra.
 Under OBC, the system supports localized edge modes with purely imaginary energies $E=\pm iu$.
@@
 3(a,b)].
-Using the standard correlation–entanglement mapping $\nu=(1+e^{\epsilon})^{-1}$, equivalently
+Using the standard correlation–entanglement mapping $\nu=(1+e^{\epsilon})^{-1}$ [36], equivalently
 $\epsilon=\log\frac{1-\nu}{\nu}=\pm 2i\arctan(2\mathcal{I}),$ (6)
 corresponding to a purely imaginary entanglement energy.
-This generalizes the Li–Haldane picture: although both energy are no longer real or degenerate, the one-to-one counting and the vanishing real parts ($\operatorname{Re}[E]=\operatorname{Re}[\epsilon]=0$) still match.
+This generalizes the Li–Haldane picture [134]: although both energy are no longer real or degenerate, the one-to-one counting and the vanishing real parts ($\operatorname{Re}[E]=\operatorname{Re}[\epsilon]=0$) still match.
 Refer to caption(a)
@@
 In both cases, the number of complex-conjugate pairs matches the number of edge modes.
-A key consequence concerns the imaginary part of the entanglement entropy.
+A key consequence concerns the imaginary part of the bulk entanglement entropy under PBC.
 Requiring the correct c=2c=-2 scaling at all critical points forces us to adopt the same branch-cut convention used in the previous section.
@@
 This identifies a subleading imaginary contribution $i\pi\omega$.
-Unlike in Hermitian SEQC, where topological features at criticality are reflected only in the entanglement spectrum [101, 83], the nH case renders them directly visible in the imaginary part of the entanglement entropy—a feature unique to nH systems.
+Unlike in Hermitian SEQC, where topological degeneracies can be diagnosed via the Affleck–Ludwig boundary $g$ function [80] in the entanglement entropy under OBC and entanglement spectrum under PBC [105, 87, 126], the nH case renders them directly visible in the imaginary part of the bulk entanglement entropy—a feature unique to nH systems.
 It is then natural to extend the same branch choice into the adjacent SPT phase by continuity.
 With that extension, $\operatorname{Im}[S_{A}]$ takes the same constant value throughout each fixed-$\omega$ sector.
-Thus $\operatorname{Im}S_{A}$ serves as a topological index reflecting the winding number $\omega$ across the $\mathcal{PT}$-symmetric sector.
+Thus $\operatorname{Im}[S_{A}]$ serves as a topological index reflecting the winding number $\omega$ across the $\mathcal{PT}$-symmetric sector.
 $\mathcal{PT}$ symmetry-protected topological edge states at non-unitary criticality.—The $\alpha$-nH SSH chain supports gapless topological features at criticality.
@@
 Shown are the mass profile $m(x)$ (black), the spectral gap $\Delta(x)=\sqrt{m(x)^{2}-u(x)^{2}}$ (blue), and the bound-state density $|\psi(x)|^{2}$ (red) localized at the interface.
-Generalized mass inversion in $\mathcal{PT}$ symmetry-enriched non-Hermitian QCP—A distinctive feature of the $\alpha$-nH SSH model is the appearance of an intermediate $\mathcal{PT}$-broken symmetry-breaking phase absent in Hermitian free-fermion systems.
+Generalized mass inversion in $\mathcal{PT}$ symmetry-enriched non-Hermitian QCPs—A distinctive feature of the $\alpha$-nH SSH model is the appearance of an intermediate $\mathcal{PT}$-broken symmetry-breaking phase absent in Hermitian free-fermion systems [76].
 This additional phase shifts the location of the critical point and is closely tied to the persistence of edge modes across the transition.
-In the Hermitian $\alpha\geq 2$-chain, topological edge states rely on kinetic inversion with long-range hopping to maintain edge states at criticality.
+In the Hermitian $\alpha\geq 2$-chain, topological edge states rely on kinetic inversion [78] with long-range hopping to maintain edge states at criticality.
 By contrast, in the nH SSH model, the survival of edge modes follows from a simple generalization of mass inversion without requiring long-range hopping ($\alpha=1$).
@@
 where $m$ is the mass profile and $u$ the non-Hermitian strength. The eigenvalue equations
-$(-\partial_{x}+m)\psi_{B}(x)$ $=(E-iu)\psi_{A}(x),$ (9)
+$(-\partial_{x}+m)\psi_{B}(x)$ $=(E-iu)\psi_{A}(x),$
 $(\partial_{x}+m)\psi_{A}(x)$ $=(E+iu)\psi_{B}(x)$ (9)
-decouple at $E=\pm iu$, yielding localized edge solutions with decay length $\sim 1/m$ for $m>0$ even as the bulk gap $\sqrt{m^{2}-u^{2}}$ closes.
-More generally, this is due to when we have more than one mass operators $\{m_{a}\Gamma_{a}\}$ with $\Gamma_{a}$ being the gamma matrices in the Dirac-like Hamiltonian.
-Here some of the $m_{a}$ can be imaginary.
-Edge localization is set by the domain-wall mass—the component that flips across the boundary—so $\xi\sim 1/|m_{bdy}|$.
-The bulk gap, however, is controlled by the spectral mass $\Delta=\sqrt{\sum_{a}m_{a}^{2}}$.
-Thus these scales are able to decouple when we have multiple mass term.
-Including this scale decoupling together with ordinary mass inversion is what we refer to as generalized mass inversion.
-This mechanism is absent in Hermitian systems, where the gap is always enhanced and cannot close without delocalizing the edge modes.
-This makes a free-fermion SEQC without long-range hopping possible only in the nH setting.
-For higher $\alpha$, a generalized mass inversion coexists with kinetic inversion, see the SM [122].
+decouple at $E=\pm iu$, yielding localized edge solutions with decay length $\sim 1/m$ for $m>0$ even as the spectrum gap $\sqrt{m^{2}-u^{2}}$ closes.
+More generally, when more than one mass term $m_{a}$ appears in the Hamiltonian, take $M=\sum_{a}m_{a}\sigma_{a}$.
+There are two scales that control the physics here: (i) the decay rate $\kappa$, which is set by the domain-wall mass—the component that flips sign across the boundary—so $\kappa\simeq-|m_{\mathrm{bdy}}|$; (ii) the spectrum gap $\Delta$, which is set by the total mass, $\Delta=\sqrt{\sum_{a}m_{a}^{2}}$.
+Thus we can see that these two scales can decouple when there are multiple mass terms.
+We refer to this scale decoupling, together with ordinary mass inversion at the boundary, as generalized mass inversion.
+For a nH SSH model, an imaginary mass $iu$ comes in so that $\Delta=\sqrt{m^{2}-u^{2}}$.
+Tuning $u\to m$ closes the bulk gap $(\Delta\to 0)$ while keeping the domain-wall mass $m$ finite, hence realizing a SEQC.
+In contrast, in Hermitian models there is no imaginary mass, so the gap cannot close without delocalizing the edge mode; this route is therefore unique to the non-Hermitian SEQC setting, making a free-fermion SEQC without long-range hopping possible only in the nH case.
+For higher $\alpha$, multiple coefficients enter the edge equation, and the neat single-parameter inversion picture no longer applies (see the SM [133]).
 The schematic in Fig.
 4(b) depicts an interface geometry (left: trivial gapped phase; right: nH SSH at its topological QCP).
-Unlike the OBC case, the interface equations remain coupled and lack a simple closed-form solution; nevertheless, generalized mass inversion pins an edge state at the interface [122].
+Unlike the OBC case, the interface equations remain coupled; while the lattice calculation is complex, we obtain an explicit continuum (field-theory) solution demonstrating that generalized mass inversion pins an edge state at the interface [133].
 Discussion and concluding remarks.—To summarize, we uncover a new class of non-unitary criticality enriched by $\mathcal{PT}$ symmetry, which hosts robust topological edge states.
@@
 More importantly, we unambiguously demonstrate that these nH critical points support nontrivial edge states protected by $\mathcal{PT}$ symmetry.
-The associated topological degeneracy is reflected in the bulk entanglement spectrum—representing a generalized Li-Haldane correspondence at nH criticality—and, remarkably, is also encoded in the imaginary part of the entanglement entropy, a feature unique to nH systems.
+The associated topological degeneracies ars reflected in the bulk entanglement spectrum—representing a generalized Li-Haldane correspondence at nH criticality—and, remarkably, is also encoded in the imaginary part of the entanglement entropy, a feature unique to nH systems.
 Finally, we analytically show that the topological edge modes at these QCPs are stabilized by a new mechanism, generalized mass inversion—without requiring long-range hopping—highlighting a distinct protection mechanism from Hermitian counterparts.
 Looking ahead, identifying symmetry-enriched nH criticalities in higher dimensions and in interacting many-body systems represents an interesting direction for future work.
-Regarding experimental realization, we note that nH free-fermion lattice models and entanglement-based observables can be implemented and probed using quantum and classical simulators, including ultracold atoms [123, 124], superconducting quantum processors [125, 126], and phononic systems [127].
+Regarding experimental realization, non-Hermitian quantum systems can be simulated on photonic platforms, where gain and loss are experimentally implemented [135, 136, 137, 138].
+Although the main observables on these platforms typically relate to single-particle physics, the survival of topologically protected edge modes within a fine-tuned gapless spectrum can still be experimentally tested [139].
+Furthermore, entanglement properties of non-Hermitian many-body states could potentially be measured in spin systems (via a Jordan-Wigner transformation) using post-selection [140].
+This investigation, however, is a direction for future work beyond the scope of our study.
 Our work not only generalizes the notion of topology to the realm of non-unitary criticality but also contributes to a deeper understanding of critical phenomena in nH systems.
@@
 Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science, and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities.
+Supplemental Material for $\mathcal{PT}$ symmetry-enriched non-unitary criticality”
+I The relation between Berry phase and winding number in non-Hermitian systems
+In this section, we review the complex Berry phase, following Refs. [132, 28]. Recall the non-Hermitian SSH Bloch Hamiltonian
+$H(k)=\begin{pmatrix}iu&v_{k}\\ v_{k}^{*}&-iu\end{pmatrix}$ (S1)
+with $u\in\mathbb{R}$. The lower band eigenstate can be express as
+[…]
+where $\varphi(k)\equiv\arg v_{k}$ and $|]\theta(k)=\arctan[iu/\absolutevalue{v_{k}}$]. The biorthogonal Berry connection and Zak phase are
+[…]
+Substituting (S2) into (S3) yields
+$A_{--}(k)=-\frac{1}{2}\partial_{k}\varphi(k)+\frac{1}{2}\partial_{k}\varphi(k)\cos\theta(k).$ (S4)
+Using $\cos\theta(k)=iu/\sqrt{|v_{k}|^{2}-u^{2}}$ we obtain
+$A_{--}(k)=-\frac{1}{2}\partial_{k}\varphi(k)+\frac{1}{2}\partial_{k}\varphi(k)\frac{iu}{\sqrt{|v_{k}|^{2}-u^{2}}}.$ (S5)
+On the $\mathcal{PT}$-symmetric part of the phase diagram—i.e., wherever the spectrum is real ($E(k)\in\mathbb{R}_{\geq 0}$, which includes the critical locus)—the factor $iu/\sqrt{|v_{k}|^{2}-u^{2}}$ is imaginary across the momentum space.
+Hence the second term of (S5) contributes only to the imaginary part, and
+$\operatorname{Re}[A_{--}](k)=-\frac{1}{2}\partial_{k}\varphi(k).$ (S6)
+Integrating over the Brillouin zone gives the quantized relation
+$\operatorname{Re}[Q_{-}]=-\frac{1}{2}\int_{-\pi}^{\pi}\partial_{k}\varphi dk=\pi\omega,\qquad\omega=\frac{1}{2\pi}\int_{-\pi}^{\pi}\partial_{k}\arg v_{k}dk\in\mathbb{Z}.$ (S7)
+Thus, throughout the $\mathcal{PT}$-symmetric part, the real Zak phase equals $\pi\omega$, and the system has protected topology.
+Inside the $\mathcal{PT}$-broken wedge there exists a nonempty set $I=\{k:\ |v_{k}|<u\}$ on which $E(k)=i\sqrt{u^{2}-|v_{k}|^{2}}$ and
+$\frac{iu}{\sqrt{|v_{k}|^{2}-u^{2}}}=\frac{u}{\sqrt{u^{2}-|v_{k}|^{2}}}\in\mathbb{R}\ (k\in I).$ (S8)
+The connection then acquires an additional real contribution on $I$,
+$\mathrm{Re}A_{--}(k)=-\frac{1}{2}\partial_{k}\varphi(k)+\frac{1}{2}\partial_{k}\varphi(k)\frac{u}{\sqrt{u^{2}-|v_{k}|^{2}}}\quad(k\in I),$ (S9)
+and the Zak phase becomes
+$\mathrm{Re}Q_{-}=\pi\omega+\frac{1}{2}\int_{I}dk\,\partial_{k}\varphi(k)\frac{u}{\sqrt{u^{2}-|v_{k}|^{2}}}.$ (S10)
+The extra integral is non-quantized, so the Berry phase no longer tracks the winding number; the winding-number (and edge-mode) count persists but is no longer robust in this regime.
+II Correlation-matrix method for quantum entanglement in non-Hermitian systems
+We consider a quadratic (free) fermion system whose ground state is Gaussian. For a bipartition $A\cup\bar{A}$, the reduced state on $A$ can be written as
+$\rho_{A}=Z^{-1}e^{-\mathcal{H}_{E}},\qquad\mathcal{H}_{E}=\sum_{i,j\in A}h^{E}_{ij}c_{i}^{\dagger}c_{j}.$ (S11)
+The correlation matrix on $A$ is defined in trace form by
+[…]
+Equivalently, as a projector onto the occupied single-particle modes,
+$C_{ij}=\sum_{\alpha\in\mathrm{occ}}L_{\alpha i}\,R^{\dagger}_{\alpha j}.$ (S13)
+Diagonalize the single-particle entanglement Hamiltonian with a similarity transform
+$h^{E}=S\mathrm{diag}(\varepsilon_{n})S^{-1}.$ (S14)
+The new basis can be expressed as $f_{L}=S\phi S^{-1}$ and $f_{R}^{\dagger}=S\phi^{\dagger}S^{-1}$, so that $\{f_{L,m},f_{R,n}^{\dagger}\}=\delta_{mn}$. In this basis
+$\mathcal{H}_{E}=\sum_{n}\varepsilon_{n}\,f_{R,n}^{\dagger}f_{L,n},\qquad\rho_{A}=\prod_{n}\frac{e^{-\varepsilon_{n}f_{R,n}^{\dagger}f_{L,n}}}{1+e^{-\varepsilon_{n}}},$ (S15)
+which gives the single-mode occupations
+$\nu_{n}\equiv\expectationvalue{f_{R,n}^{\dagger}f_{L,n}}=\frac{1}{1+e^{\varepsilon_{n}}}.$ (S16)
+The von Neumann entropy can be written as
+[…]
+=Tr[$=-\Tr\left[\frac{e^{-h^{E}}}{1+e^{-h^{E}}}\ln\frac{e^{-h^{E}}}{1+e^{-h^{E}}}+\frac{1}{1+e^{-h^{E}}}\ln\frac{1}{1+e^{-h^{E}}}\right]$
+$=-\sum_{n}\left[\nu_{n}\ln\nu_{n}+(1-\nu_{n})\ln(1-\nu_{n})\right],$ (S17)
+which is the expression used in the main text.
+III Details of the branch-cut prescription in non-Hermitian systems
+III.1 Edge–mode
+In nH SSH chains, the biorthogonal correlation spectrum of a subsystem may contain complex eigenvalues. At topological QCPs, one observes a complex-conjugate pair (see Fig. S1 (a))
+$\nu_{\pm}=\frac{1}{2}\pm i\,\mathcal{I}$$\qquad\mathcal{I}\in\mathbb{R_{+}},$ (S18)
+which contributes to the von Neumann entropy
+$S_{A}=-\sum_{n}\Big[\nu_{n}\log\nu_{n}+(1-\nu_{n})\log(1-\nu_{n})\Big].$ (S19)
+Note that the logarithm here is a multivalued function: for any nonzero $z=|z|e^{i\phi}$,
+$\log z=\ln|z|+i\big(\phi+2\pi m\big),\qquad m\in\mathbb{Z},$ (S20)
+which gives rise to a branch-choice ambiguity in $S_{A}$ that we resolve in this work.
+As stated in the main text, this complex-conjugate pair corresponds to an edge mode. We denote its contribution by
+[…]
+On the principal branch, with $\mathrm{Arg}\in(-\pi,\pi]$, write $\frac{1}{2}\pm i\mathcal{I}=r\,e^{\pm i\phi}$, where $r=\sqrt{\frac{1}{4}+\mathcal{I}^{2}}$ and $\phi=\arctan(2\mathcal{I})\in(0,\pi)$.
+Substituting into (S21) yields
+$S_{\rm edge}^{(\rm princ)}=-2\ln r+4\phi\mathcal{I},$ (S22)
+which is purely real. However, this choice leads to scaling behavior that deviates markedly from the expected conformal form (see Fig. S1 (b)), as pointed out in previous work [36].
+From Fig.
+S1 (b), the principal-branch fit yields an entanglement entropy that is positive and increases with subsystem size $\ell_{A}$, in tension with the expected conformal scaling.
+Inspecting Eq.
+(S22), we single out the term $4\phi\mathcal{I}$ as a candidate culprit for the positive deviation.
+At criticality the edge-pair parameter $\mathcal{I}$ is parametrically large (formally $|\mathcal{I}|\to\infty$ if the EP is exactly occupied), so with $\phi=\arctan(2\mathcal{I})$ we have
+$\phi\to\frac{\pi}{2}\qquad\Rightarrow\qquad 4\phi\mathcal{I}\longrightarrow 2\pi\mathcal{I}.$ (S23)
+This suggestive form motivates a minimal fix.
+The factor $4\phi$ in front of $\mathcal{I}$ originates from the phase choices of the four logarithmic terms.
+By shifting the branch of one of these logarithms by $2\pi$ in its argument, we replace $4\phi\mathcal{I}$ with $(4\phi-2\pi)\mathcal{I}$, thereby canceling this positive contribution.
+With this branch choice, the edge contribution becomes
+$S_{\rm edge}^{(\rm branch)}=-2\ln r+\big(4\phi-2\pi\big)\mathcal{I}\pm i\pi.$ (S24)
+We fix the sign convention to the negative one, yielding $\operatorname{Im}[S_{\rm edge}]=-\pi$. Consequently, the entanglement entropy reads
+$S_{A}=-\frac{2}{3}\log\ell_{A}-i\pi+\text{const.}.$ (S25)
+As shown in Fig. S1 (c), with this branch reassignment in place, the real-part fits recover the expected conformal scaling. Thus we regard this as the correct, physically meaningful branch choice.
+Refer to caption(a)
+Refer to caption(b)
+Refer to caption(c)
+Figure S1: Comparison of branch choices for entanglement in the non-Hermitian SSH chain at its topological QCP ($L=10^{4}$; $(v,w,u)=(1,2,1-10^{-12})$).
+(a) Entanglement spectrum $\{\nu\}$ at $\ell_{A}=100$ (left: $\operatorname{Re}[\nu]$; right: $\operatorname{Im}[\nu]$).
+(b) $S_{A}(\ell_{A})$ evaluated with the principal branch.
+(c) $S_{A}(\ell_{A})$ on the physically consistent branch, exhibiting the expected c=2c=-2 scaling.
+Blue (red) markers denote $\operatorname{Re}[S_{A}]$ ($\operatorname{Im}[S_{A}]$).
+III.2 Gapped mode
+Prior sections clarified that the pathology in the entanglement entropy of the topological nH SSH model is a branch-choice artifact.
+Earlier studies [36] also reported an "entropy flip" at small gaps or under slight boundary deviations, attributing it to boundary or exceptional-point (EP) sensitivity.
+Re-examining the gapped regime, we find that the apparent entropy flip coincides with the onset of a complex quartet in the correlation spectrum, which we refer to as gapped modes
+$\{\nu,\;\nu^{*},\;1-\nu,\;1-\nu^{*}\}$ (S26)
+in the correlation spectrum. This indicates that the discontinuity is not a genuine phase transition but another branch-choice artifact.
+The quartet contribution to the entanglement entropy is
+$S_{\rm quartet}=-2\Big[\nu\log\nu+(1-\nu^{*})\log(1-\nu^{*})\Big]+\text{c.c.},$ (S27)
+where $\nu=\mathcal{R}+i\,\mathcal{I}$ with $\mathcal{I}\in\mathbb{R_{+}}$. On the principal branch, write
+$\nu=re^{i\phi},\qquad 1-\nu^{*}=\rho e^{i\varphi},\qquad r=|\nu|,\rho=|1-\nu|,\phi=\arg(\nu),\varphi=\arg(1-\nu^{*}),\arg\in(-\pi,\pi].$ (S28)
+Substituting into (S27) gives
+$S_{\rm quartet}^{(\rm princ)}=-4\mathcal{R}\ln r-4(1-\mathcal{R})\ln\rho+4\mathcal{I}(\phi+\varphi),$ (S29)
+which is purely real on the principal branch.
+However, as noted above, numerics show an apparent "flip" in the entropy precisely when a complex quartet appears in the spectrum (see Fig.
+S2 (a-c)), signaling a branch-choice artifact addressed below.
+To resolve this branch-choice artifact, we adopt the same strategy used for the edge pair.
+The difference here is that a quartet contributes eight phases in total.
+Shifting the branches of two of these logarithms by $2\pi$ yields an expression analogous to Eq.
+(S24).
+In the quartet case the imaginary parts cancel exactly between conjugate contributions, and we obtain
+$S_{\rm quartet}^{(\rm branch)}=-4\mathcal{R}\ln r-4(1-\mathcal{R})\ln\rho+4\mathcal{I}\big(\phi+\varphi-\pi\big),$ (S30)
+which is purely real. With this branch assignment, the entropy recovers the smooth, gapped behavior expected in this regime (see Fig. S2 (d-e)).
+Refer to caption(a)
+Refer to caption(b)
+Refer to caption(c)
+Refer to caption(d)
+Refer to caption(e)
+Figure S2: Comparison of branch choices for entanglement in a small-gap non-Hermitian SSH chain ($L=10^{4}$; $(v,w,u)=(2,1,1-10^{-7})$).
+(a–b) Entanglement spectrum $\{\nu\}$ around the onset: a complex quartet first appears at $\ell_{A}=11$ (absent at $\ell_{A}=10$).
+(c) Principal-branch $S_{A}(\ell_{A})$ shows an apparent "flip" at $\ell_{A}=11$, coincident with the quartet onset.
+(d) Using a physically consistent branch removes the flip; for small gaps the scaling is consistent with c=2c=-2.
+(e) Correct-branch $S_{A}(\ell_{A})$ for $u=1-10^{-7},\,1-10^{-5},\,1-10^{-3},\,1-10^{-1}$: the curves flatten smoothly with increasing gap $\Delta$ and approach the trivial entanglement (horizontal line).
+III.3 Symmetry constraints: why symmetric pairs and quartets appear
+Having shown that the entanglement entropy is well-defined whenever the correlation spectrum is symmetric, we now recall why this symmetry holds: microscopic symmetries enforce a pairing structure in the correlation spectrum [142].
+We write the Bloch Hamiltonian in the minimal $\mathcal{PT}$-symmetric form
+$H(k)=d_{x}(k)\sigma_{x}+d_{y}(k)\sigma_{y}+iu(k)\sigma_{z},\qquad d_{x},d_{y},u\in\mathbb{R}.$ (S31)
+There are two symmetries:
+(i) time-reversal–like symmetry ($T_{+}$):
+$H(k)^{*}=d_{x}\sigma_{x}-d_{y}\sigma_{y}-iu\sigma_{z},$ (S32)
+and conjugation by $\sigma_{x}$ gives
+$\sigma_{x}H(k)^{*}\sigma_{x}=d_{x}\sigma_{x}+d_{y}\sigma_{y}+iu\sigma_{z}=H(k).$ (S33)
+(ii) pseudo-Hermiticity ($PH_{-}$):
+$H(k)^{\dagger}=d_{x}\sigma_{x}+d_{y}\sigma_{y}-iu\sigma_{z},$ (S34)
+and conjugation by $\sigma_{z}$ yields
+$\sigma_{z}H(k)^{\dagger}\sigma_{z}=-d_{x}\sigma_{x}-d_{y}\sigma_{y}-iu\sigma_{z}=-H(k).$ (S35)
+These two Hamiltonian symmetries imply two constraints on the correlation spectrum, yielding the pairing structures used above.
+[…]
+[…]
+together with $T_{+}cT_{+}^{-1}=U_{T}c$ and $T_{+}c^{\dagger}T_{+}^{-1}=c^{\dagger}U_{T}^{\dagger}$, we obtain
+[…]
+[…]
+If the state is also invariant under time-reversal-like symmetry, we have
+$C=U_{T}C^{*}U_{T}^{\dagger}.$ (S38)
+The same relation holds for the subsystem matrix $C_{A}$ because $U_{T}$ acts on-site and thus commutes with the spatial restriction:
+$C_{A}=U_{T}C_{A}^{*}U_{T}^{\dagger}.$ (S39)
+Consequently, the spectrum is closed under complex conjugation:
+$\nu\leftrightarrow\nu^{*}.$ (S40)
+Pseudo-Hermiticity ($PH_{-}$).
+Using Eq.
+(S35), $u_{\mathrm{ph}}H^{\dagger}u_{\mathrm{ph}}^{\dagger}=-H$ (with $u_{\mathrm{ph}}=\sigma_{z}$ in the Bloch basis), we now derive the corresponding constraint on the correlation matrix and the induced pairing of its eigenvalues.
+[…]
+[…]
+[…]
+[…]
+for some phases $e^{i\alpha_{n}}$ and nonzero normalizations $\mathcal{N}_{n}$ (fixed by gauge). The correlation operator can be written as
+$|,C=\sum_{n}s_{n}\ket{R_{n}}\bra{L_{n}}$, (S43)
+with complex occupancies $s_{n}$; at biorthogonal half filling these satisfy $s_{n}^{*}+s_{-n^{*}}=1$ pairwise.
+Conjugating $C^{\dagger}$ by $u_{\mathrm{ph}}$ and using the relations above,
+$|.u_{\mathrm{ph}}C^{\dagger}u_{\mathrm{ph}}^{\dagger}=\sum_{n}s_{n}^{*}\ket{R_{-n^{*}}}\bra{L_{-n^{*}}}$. (S44)
+Using $s_{n}^{*}+s_{-n^{*}}=1$ and completeness then gives the compact identity
+$u_{\mathrm{ph}}C^{\dagger}u_{\mathrm{ph}}^{\dagger}+C=\mathbf{I}.$ (S45)
+Because $u_{\mathrm{ph}}$ acts on-site, the same holds after restricting to the subsystem $A$:
+$u_{\mathrm{ph}}C_{A}^{\dagger}u_{\mathrm{ph}}^{\dagger}+C_{A};=\mathbf{I}.$ (S46)
+Thus the spectrum is closed under the $PH_{-}$ mapping:
+$\nu\leftrightarrow 1-\nu^{*}.$ (S47)
+Together with (S40) ($\nu\leftrightarrow\nu^{*}$), this ensures that every complex eigenvalue groups into the required pairs or fourfold sets, underpinning our branch-cut prescription.
+III.4 Comparison with the absolute-value prescription
+An alternative prescription from prior work [55] reproduces the central charge at the nH SSH topological QCP by replacing the multivalued complex logarithm in the entropy with $|\log\absolutevalue{\,\cdot\,}$,
+$S_{A}^{(\mathrm{abs})}=-\sum_{n}\big[\nu_{n}\log|\nu_{n}|+(1-\nu_{n})\log|1-\nu_{n}|\big],$ (S48)
+i.e. effectively $S=-\mathrm{Tr}\rho\log\rho\mapsto-\mathrm{Tr}\rho\log|\rho|$ at the correlation-spectrum level. For the edge pair $\nu_{\pm}=\tfrac{1}{2}\pm i\mathcal{I}$ this gives
+$S_{\rm edge}^{(\mathrm{abs})}=-2\ln r,\qquad r=\sqrt{\tfrac{1}{4}+\mathcal{I}^{2}},$ (S49)
+whereas our branch-cut prescription yields
+$S_{\rm edge}^{(\mathrm{branch})}=-2\ln r+(4\phi-2\pi)\mathcal{I}-i\pi,\qquad\phi=\arctan(2\mathcal{I}).$ (S50)
+Thus the absolute-value method removes the $4\phi\mathcal{I}$ term by construction. To compare the two methods, we therefore examine the real part of $S_{A}$, the difference is
+[…]
+[…]
+$S_{\rm quartet}^{(\mathrm{abs})}=-4\mathcal{R}\ln r-4(1-\mathcal{R})\ln\rho,\qquad S_{\rm quartet}^{(\mathrm{branch})}=S_{\rm quartet}^{(\mathrm{abs})}+4\mathcal{I}(\phi+\varphi-\pi),$ (S52)
+with $r=|\nu|$, $\rho=|1-\nu|$, $\phi=\arg\nu$, and $\varphi=\arg(1-\nu^{*})$.
+The discrepancy stems from this quartet contribution, so a clear difference appears as soon as a quartet first enters the entanglement spectrum.
+At that point the absolute-value prescription typically shows a kink, whereas the branch-cut prescription remains smooth (see Fig.
+S3(a-b)).
+[…]
+In summary, both prescriptions reproduce the real scaling at the topological QCP of the ordinary nH–SSH chain (up to an irrelevant constant offset), but the branch-cut prescription preserves the original definition $S=-\mathrm{Tr}\,\rho\log\rho$ and is less sensitive to small gaps.
+Moreover, in higher-$\alpha$ models with non–mass-inversion edge modes—where the edge-mode imaginary part saturates to $\mathcal{O}(1)$—the absolute-value prescription introduces non-smooth behavior in the second derivative around this crossover.
+By contrast, the branch-cut result remains smooth.
+Taken together, the branch-cut prescription is the more convincing and physically grounded method for computing the nH entanglement entropy.
+Refer to caption(a)
+Refer to caption(b)
+Refer to caption(c)
+Refer to caption(d)
+Refer to caption(e)
+Figure S3: Comparison of entanglement–entropy prescriptions: branch-cut vs absolute value.
+Red $\times$ denotes $\mathrm{Re}[S_{A}^{(\mathrm{abs})}]$, and blue $\circ$ denotes $\mathrm{Re}[S_{A}^{(\mathrm{branch})}]$.
+Small-gap nH SSH, $(L=10000;(v,w,u)=(2,1,1-10^{-7}))$.
+A complex quartet first appears at $\ell_{A}=11$ (see Fig.
+S2(a-b)).
+(a) $\mathrm{Re}[S_{A}]$: the absolute-value prescription $\mathrm{Re}[S_{A}^{(\mathrm{abs})}]$ shows a kink at $\ell_{A}=11$, whereas the branch-cut result $\mathrm{Re}[S_{A}^{(\mathrm{branch})}]$ remains smooth.
+(b) Slope $S_{A}^{\prime}$: a clear jump occurs only for the absolute-value curve.
+Higher-$\alpha$-nH SSH $(L=100;\alpha=2;(v,w,u)=(3,1,2-10^{-15}))$.
+(c) Edge-mode imaginary part $\mathcal{I}(\ell_{A})$ drops rapidly and becomes $\mathcal{O}(1)$ for $\ell_{A}\geq 8$.
+(d) $\mathrm{Re}[S_{A}]$: once $\mathcal{I}$ is $\mathcal{O}(1)$, the two prescriptions diverge smoothly; for visual alignment we subtract the constant offset from the absolute-value curve.
+(e) Second derivative $S_{A}^{\prime\prime}$: a spike at $\ell_{A}=8$ is present for the absolute-value prescription, whereas the branch-cut result remains smooth.
+III.5 OBC/disorder: symmetry breaking and a regularized entropy
+With the branch prescription, entanglement entropy in non-Hermitian systems appears well-behaved, much as one ordinarily sees in Hermitian systems.
+However, that apparent simplicity comes with a caveat: the prescription relies not only on the microscopic symmetries of the Hamiltonian but also on the state to preserve them.
+In clean systems with periodic boundaries, both symmetries are respected by the ground state, and the correlation spectrum naturally organizes into the pairings required for a well-behaved entropy.
+Beyond this idealized setting, perturbations common in CFT studies—such as open boundary conditions and spatial inhomogeneity—typically break these symmetries at the level of the state, even when the single-particle spectrum remains real ($\mathcal{PT}$-unbroken).
+[…]
+Note that if the state preserved $\mathcal{PT}$ with on-site parity $\mathcal{P}=\sigma_{x}$, it would obey the local constraint1
+$n_{i,A}=n_{i,B}^{*}\qquad\Longleftrightarrow\qquad\operatorname{Im}[n_{i}]=0\ \text{for all }i,$ (S53)
+Under OBC, however, nonreciprocal hopping skews the biorthogonal amplitudes and introduces an imaginary tail at the boundaries.
+This boundary-localized imaginary density is direct evidence that the state breaks $\mathcal{PT}$ symmetry under OBC.
+Consequently, the conjugation pairing $\nu\leftrightarrow\nu^{*}$ is no longer guaranteed, and the symmetry of the correlation spectrum reduces to the residual $PH_{-}$ pairing $\nu\leftrightarrow 1-\nu^{*}$.
+Similar effect occurs in the presence of spatial inhomogeneity: generic perturbations induce a nonzero net imaginary amplitude throughout the bulk, again reflecting symmetry breaking at the level of the state.
+1 Why can a local particle number be complex here?
+In a $\mathcal{PT}$-symmetric nH system, the analogue of a Hermitian observable is a $\mathcal{PT}$-symmetric operator, which has a real biorthogonal expectation value by the identity (S36).
+The on-site densities $n_{i,A}$ and $n_{i,B}$ are exchanged by $\mathcal{P}$ and are therefore not $\mathcal{PT}$-symmetric; the $\mathcal{PT}$-symmetric unit-cell density $n_{i}\equiv n_{i,A}+n_{i,B}$ is the appropriate physical number to have real expectation value.
+[…]
+$\operatorname{Im}[S_{\text{gap}}]=\pm 2\pi\mathcal{R}\quad\text{or}\quad\pm 2\pi(1-\mathcal{R}),$ (S54)
+depending on which phase is shifted.
+Thus, for gapped modes the real part remains regular, but a nonvanishing, nonuniversal (branch-dependent) imaginary remainder persists and spoils the quantization of the imaginary entanglement entropy once $T_{+}$ is broken.
+The situation is more serious when edge modes are present.
+Once $T_{+}$ is broken, the edge modes no longer form the conjugate pair $\tfrac{1}{2}\pm i\mathcal{I}$; instead they take the form $\nu_{\alpha}=\tfrac{1}{2}+i\mathcal{I}_{\alpha}$ with independent $\mathcal{I}_{\alpha}$.
+Although their total number still matches the winding number and appear in pair, the pairwise structure of the imaginary parts is lost.
+This breaks the analysis in Sec.
+III.1; thus, there is no branch choice that yields a regular value and the conformal scaling.
+Although it may be acceptable that a symmetry-broken non-Hermitian system does not admit a regular entanglement entropy, we can still extend its usefulness by a minimal redefinition that restores the missing symmetry partner.
+Since the pathology arises from the loss of the conjugation partner $\nu^{*}$, a simple and operational fix is to add the complex-conjugate contribution by hand and divide by two:
+$S_{A}^{\rm reg}\equiv-\frac{1}{2}\sum_{\nu\in{\rm spec}(C_{A})}\Big\{\underbrace{\nu\log\nu+(1-\nu)\log(1-\nu)}_{\displaystyle S_{\nu}}+\underbrace{\nu^{*}\log\nu^{*}+(1-\nu^{*})\log(1-\nu^{*})}_{\displaystyle S_{\nu^{*}}}\Big\},$ (S55)
+or, in a density matrix form,
+[…]
+With this regularized entropy, OBC or disordered systems regain well-behaved scaling: the conformal scaling is recovered, and the quantized imaginary entropy is retained.
+For topological OBC chains, where the edge modes carry purely imaginary energies, the same prescription can still be used by fixing the edge filling.
+In particular, filling the two edge states as a Bell pair restores the half-filled structure (see Fig.
+S4(c)) and thereby the residual $PH_{-}$ symmetry in the entanglement spectrum.
+With this symmetric edge filling, the entanglement can be computed in a well-behaved way using the regularized entropy $S^{\rm reg}$.
+The imaginary tail in the topological QCP actually carries additional physics.
+By computing the unit cell density at the topological QCP, we find that relative to the trivial QCP, where the imaginary density accumulates at the right edge, the topological QCP shows the opposite profile [Fig.
+S4(d)].
+This sign reversal reflects which non-reciprocal hopping dominates.
+Upon tuning toward $v/w\approx 1$ (equivalently $u/w\approx 0$), the transition between two boundaries proceeds via delocalization of the imaginary density into the bulk [Fig.
+S4(e)]: the tail spreads across the chain and then relocalizes at the opposite boundary.
+Refer to caption(a)
+Refer to caption(b)
+Refer to caption(c)
+Refer to caption(c)
+Refer to caption(c)
+Figure S4: Site and unit-cell densities in the nH–SSH chain under OBC.
+(a) Particle density $n_{i}=\langle c_{i}^{\dagger}c_{i}\rangle$ for $L=100$, $(v,w,u)=(2,1,1)$.
+$\operatorname{Re}[n_{i}]$ is at half filling, while $\operatorname{Im}[n_{i}]$ reveals sublattice asymmetry that violates $\mathcal{PT}$-symmetry.
+(b) Imaginary part of the unit-cell density $n_{i}=n_{i,A}+n_{i,B}$ for the same parameters.
+$\operatorname{Im}[n_{i}]$ is nonzero near the edges, explicitly showing $\mathcal{PT}$-symmetry breaking at the boundaries.
+(c) Particle density at the topological QCP, $L=100$, $(v,w,u)=(1,2,1)$.
+A similar profile is observed; the real part returns to half filling when the edge mode is filled symmetrically.
+(d) Imaginary unit-cell density corresponding to (c): the imaginary density is localized at the opposite boundary compared to the trivial case, reflecting the reversal of the dominant nonreciprocal direction.
+(e) Imaginary unit-cell density for $(v,w,u)=(9,10,1)$ (i.e., $v/w\approx 1$): the boundary density delocalizes across the bulk during the crossover, signaling the migration of the boundary profile between edges.
+IV Scaling of entanglement entropy and ground-state Energy in higher-$\alpha$ non-Hermitian QCPs
+IV.1 PBC entanglement at higher $\alpha$
+We extend the periodic–boundary entanglement check of the central charge to higher-$\alpha$ cases not covered in the main text.
+Throughout this subsection we set $L=10^{4}$.
+For lower-winding critical points we use $(v,w,u)=(2,1,1-10^{-12})$, while for higher-winding critical points we use $(1,2,1-10^{-12})$.
+The real part of the entropy is fitted to the Calabrese-Cardy form
+$\operatorname{Re}[S_{A}^{\rm PBC}](\ell_{A})=\frac{c}{3}\log\left[\sin\left(\frac{\pi\ell_{A}}{L}\right)\right]+s_{0}$ (S57)
+The imaginary part under PBC is step-like in $\ell_{A}$: it drops by $\pi$ each time an edge mode is included and then saturates once all $\omega$ modes are covered.
+Eventually, $\operatorname{Im}S_{A}^{\rm PBC}=-\pi\omega$ for a QCP with winding number $\omega$.
+In the small-$\ell_{A}$ regime we observe a pronounced UV deviation.
+For lower-winding QCPs, this originates from unit cell granularity: a single unit cell captures at most one edge mode, so one must have $\ell_{A}>\omega$ to include all $\omega$ modes; therefore it is not surprising to see the UV deviation before this threshold.
+Furthermore, at the higher-winding QCPs, we have an additional edge mode generated by generalized mass inversion, thus with a finite decay length $\xi$, so the overlap with the cut decays smoothly over a scale $\sim\xi$; correspondingly, the deviation extends over a longer range.
+To obtain a clean Calabrese–Cardy fit, we exclude a minimal subset of contaminated UV points: for lower-winding QCPs we remove the first $N_{\rm edge}$ data points, while for higher-winding QCPs we choose the smallest trim such that the sum of squared residuals satisfies $\mathrm{SSE}\leq 10^{-4}$.
+As shown in Fig.
+S5, across all higher-$\alpha$ cases the real part of the PBC entanglement follows the Calabrese–Cardy form with slope $c/3\simeq-2/3$ once a minimal UV trim is applied.
+The imaginary part exhibits the expected step-to-saturation behavior and ultimately equals $\pi\omega$.
+Taken together, these checks confirm c=2c=-2 for the higher-$\alpha$ critical points considered here.
+Refer to caption(a)
+Refer to caption(b)
+Refer to caption(c)
+Refer to caption(d)
+Refer to caption(e)
+Refer to caption(f)
+Figure S5: PBC entanglement entropy at higher-$\alpha$, $L=10000$.
+Blue dots: $\operatorname{Re}[S_{A}]$; red dots: $\operatorname{Im}[S_{A}]$.
+Lower-winding QCPs use $(v,w,u)=(2,1,1-10^{-12})$; higher-winding QCPs use $(1,2,1-10^{-12})$.
+(a) $\alpha=2$ (lower): c/3=0.6638c/3=-0.6638.
+(b) $\alpha=3$ (lower): c/3=0.6644c/3=-0.6644.
+(c) $\alpha=3$ (higher): trim $\ell_{A}\leq 10$ (SSE $\leq 10^{-4}$); c/3=0.6668c/3=-0.6668.
+(d) $\alpha=4$ (lower): c/3=0.6658c/3=-0.6658.
+(e) $\alpha=4$ (higher): trim $\ell_{A}\leq 14$ (SSE $\leq 10^{-4}$); c/3=0.6677c/3=-0.6677.
+(f) $\alpha=5$ (lower): c/3=0.6683c/3=-0.6683.
+In all panels, $\operatorname{Im}[S_{A}]$ is step-like at small $\ell_{A}$ and saturates to $\pi\omega$ once all edge modes are covered, while $\operatorname{Re}[S_{A}]$ follows the Calabrese–Cardy scaling consistent with c=2c=-2.
+IV.2 OBC Casimir energy
+We begin from the standard Casimir forms for open and periodic chains,
+$E_{0}^{\rm OBC}(L)=L\,\epsilon+b-\frac{\pi v_{F}c}{24\,L}+\mathcal{O}(L^{-2}),\qquad E_{0}^{\rm PBC}(L)=L\,\epsilon-\frac{\pi v_{F}c}{6\,L}+\mathcal{O}(L^{-2}),$ (S58)
+where $\epsilon$ is the bulk energy density, $b$ the order-$L^{0}$ boundary constant, and $v_{F}$ the bulk velocity (fixed to $v_{F}=\sqrt{2}$).
+Following boundary CFT and surface critical analyses, leading boundary corrections can be absorbed into an extrapolation length—a microscopic shift of the effective boundary position—implemented by replacing $L\to L_{\rm eff}=L+\Delta_{L}$ in the Casimir term [143, Stéphan_2013].
+This yields the working-fit form,
+$E_{0}^{\rm OBC}(L)=L\,\epsilon+b-\frac{\pi v_{F}c}{24\,(L+\Delta_{L})}\,$, (S59)
+with a non-universal $\Delta_{L}=\mathcal{O}(1)$ that depends on the boundary class.
+Using Eq. (S59), the expected $1/L$ slopes follow immediately. For c=2c=-2 and $v_{F}=\sqrt{2}$,
+$A_{\rm OBC}=\frac{\pi v_{F}(-c)}{24}\approx 0.370240,\qquad A_{\rm PBC}=\frac{\pi v_{F}(-c)}{6}\approx 1.480961.$ (S60)
+For reference, the PBC fit reported in the main text [Fig.
+2(d)] yields $A_{\rm PBC}=1.4807$, which is in excellent agreement with c=2c=-2.
+A single non-universal, $\mathcal{O}(1)$ extrapolation length per boundary class brings the OBC data onto this line:
+$\alpha=1:\ \Delta_{L}^{\rm triv}=+2,\ \ \Delta_{L}^{\rm topo}=-1;\qquad\alpha=2:\ \Delta_{L}^{\rm lower}=+1,\ \ \Delta_{L}^{\rm higher}=-2.$ (S61)
+With these shifts, the fits in Fig. S6 align with the c=2c=-2 prediction.
+Refer to caption(a)
+Refer to caption(b)
+Refer to caption(c)
+Refer to caption(d)
+Figure S6: OBC Casimir scaling.
+Numerical fit values (slope AA of $E_{0}^{\rm OBC}$ vs $1/L_{\rm eff}$ and constant $b$): (a) $\alpha=1$ trivial QCP: $A=0.3701466$, b=0.0472026b=-0.0472026; (b) $\alpha=1$ topological QCP: $A=0.3702412$, $b=1.0749959$; (c) $\alpha=2$ lower winding QCP: $A=0.3701497$, $b=1.753430053$; (d) $\alpha=2$ higher winding QCP: $A=0.3702569$, $b=2.8756283$.
+All slopes are consistent with the c=2c=-2 expectation for $v_{F}=\sqrt{2}$.
+Beyond the $1/L$ term, the order-$L^{0}$ constant $b$ encodes the contibution of boundary entropy (Affleck–Ludwig) $\ln g$.
+While absolute values are scheme dependent, differences at a fixed UV scheme are meaningful.
+In Fig.
+S6, the lower-winding baseline at $\alpha=1$ is close to zero, and—within a given $\alpha$—the higher–lower winding QCP offset in $b$ is of the order $\ln 2$, consistent with an additional edge mode.
+Beyond the c=2c=-2 Casimir term, boundary entropy differences encode topology and reveal symmetry-enriched criticality [80].
+IV.3 OBC entanglement
+We now extend the entanglement check to OBC with the regularized entropy defined in Sec.
+III.5.
+In this setting, we observe a comparable length offset from UV physics; unlike the Casimir case, where the extrapolation length is pinned to integer values, the OBC offset is parameter-dependent and continuous.
+To preserve the empirical minimum at half-chain we impose
+$\Delta L=2\,\Delta\ell,$ (S62)
+and fit the real part with
+$\operatorname{Re}S_{A}^{\rm OBC}(\ell_{A})=\frac{c}{6}\log\!\left[\sin\!\left(\frac{\pi[\ell_{A}+2\Delta\ell]}{L+2\Delta\ell}\right)\right]+s_{0}^{\prime}.$ (S63)
+Notice that the shift $\Delta L$ here is different from the Casimir shift.
+This additional term is not an ad-hoc adjustment.
+In the long-wavelength limit the $\mathcal{O}(1)$ offset becomes negligible and the original Calabrese–Cardy form yields a good fit.
+Introducing $\Delta L=2\Delta\ell$ simply accelerates convergence at accessible sizes by removing a model-dependent UV offset; it does not alter the universal coefficient $c/6$.
+We can investigate the origin of this UV offset by varying $v/w$ at trivial QCP of the nH SSH chain under OBC.
+In Fig.
+S7(a), a case with $v/w$ of order one exhibits only a small shift.
+In contrast, Fig.
+S7(b) presents $v/w\approx 1$ (equivalently $u/v\approx 0$), where the shift becomes large and a pronounced UV offset appears.
+Recalling the discussion of the imaginary tail in Sec.
+III.5, these trends indicate that the UV offset tracks the tail’s localization length: when the tail is tightly localized the offset is small, whereas near $v/w\approx 1$ the tail penetrates deeper into the bulk and the offset grows accordingly.
+This explains why a nonzero offset continuously arises even at the trivial QCP without an edge mode, and why its magnitude depends smoothly on the coefficients.
+In both panels, the fitted slope coincides with $c/6$ (for c=2c=-2).
+Then we investigate the UV offset dependence on $v/w$ at topological QCP.
+In Fig.
+S7(c) (the $\alpha=1$ topological QCP with $v/w=\mathcal{O}(1)$) the required shift is negative, in line with the effective-length reduction also seen in the Casimir analysis for topological boundaries.
+Moving toward $v/w\approx 1$ [Fig.
+S7(d), equivalently $u/w\approx 0$], the offset increases markedly in magnitude, consistent with a slowly decaying higher-winding imaginary tail sampled from the boundary side; the entanglement edge mode becomes visible only at larger subsystem sizes (around $\ell_{A}\approx 14$), so the imaginary part relaxes over a long distance.
+In both panels, the fitted slope coincides with $c/6$ (for c=2c=-2).
+Having verified the shifted Calabrese–Cardy formula in Eq.
+(S63) on the $\alpha=1$ nH SSH cases, we apply the same analysis to $\alpha=2$ in Fig.
+S7(e–f).
+In both lower and higher winding QCPs, a single boundary-dependent $\Delta\ell$ collapses $\operatorname{Re}S_{A}$ onto the Calabrese–Cardy form, while $\operatorname{Im}S_{A}$ exhibits the expected step-to-plateau behavior.
+In all panels, the fitted slope coincides with $c/6$ (for c=2c=-2).
+In summary, despite UV offsets (continuous $\Delta\ell$ shifts and minimal trims), the entanglement under OBC in the $\alpha$-nH SSH models follows the Calabrese–Cardy form with slope $c/6$, consistently confirming c=2c=-2 across all critical points considered.
+Refer to caption(a)
+Refer to caption(b)
+Refer to caption(c)
+Refer to caption(d)
+Refer to caption(e)
+Refer to caption(f)
+Figure S7: OBC entanglement entropy $S_{A}$ as a function of subsystem size $\ell_{A}$ for $L=200$.
+Blue dots show $\operatorname{Re}S_{A}$, red dots show $\operatorname{Im}S_{A}$; solid blue curves are the numerical fits.
+For each panel we exclude the smallest-$\ell_{A}$ data until $\mathrm{RMSE}\leq 10^{-4}$, then fit $c/6$, $s_{0}$, and $\Delta\ell$.
+(a) $(v,w,u)=(2,1,1)$, $\alpha=1$: c/6=0.3344c/6=-0.3344, $s_{0}=-1.2128$, $\Delta\ell=1.1048$, trim $\ell_{A}\leq 2$.
+(b) $(10,9,1)$, $\alpha=1$: c/6=0.3368c/6=-0.3368, $s_{0}=-0.4212$, $\Delta\ell=5.4781$, trim $\ell_{A}\leq 7$.
+(c) $(1,2,1)$, $\alpha=1$: c/6=0.3321c/6=-0.3321, $s_{0}=-0.8610$, $\Delta\ell=-0.6154$, trim $\ell_{A}\leq 2$.
+(d) $(9,10,1)$, $\alpha=1$: c/6=0.3377c/6=-0.3377, $s_{0}=0.2512$, $\Delta\ell=-4.1113$, trim $\ell_{A}\leq 21$.
+(e) $(2,1,1)$, $\alpha=2$: c/6=0.3348c/6=-0.3348, $s_{0}=-0.5719$, $\Delta\ell=0.6487$, trim $\ell_{A}\leq 3$.
+(f) $(1,2,1)$, $\alpha=2$: c/6=0.3303c/6=-0.3303, $s_{0}=-0.4067$, $\Delta\ell=-1.4857$, trim $\ell_{A}\leq 8$.
+The UV offset $\Delta\ell$ is governed by the ratio $v/w$ (larger as $v/w\to 1$), while the trim distance reflects the presence of a mass-inversion edge mode: only the higher-winding QCP requires a large trim window.
+Across all panels the fits reproduce the Calabrese–Cardy slope $c/6$, consistent with c=2c=-2.
+V The physical mechanism of generating topological edge states at higher-$\alpha$ non-Hermitian criticality
+We start from the block Hamiltonian of the $\alpha$-non-Hermitian SSH model
+$\mathcal{H}(k)=\begin{bmatrix}iu&v_{k}\\ v_{k}^{*}&-iu\end{bmatrix},\qquad v_{k}=v\,e^{-i(\alpha-1)k}-w\,e^{-i\alpha k},$ (S64)
+and pass to a continuum description via the non-local map
+$e^{ik}\ \Rightarrow\ (\partial_{x}+1).$ (S65)
+We use this replacement—rather than expanding $v_{k}$ to $\alpha$-th order in kk—because finite Taylor series mix different hopping ranges into the same derivative coefficients, letting near-neighbor terms renormalize those that should represent longer-range physics.
+In contrast, the map $e^{-ink}\mapsto(\partial_{x}+1)^{n}$ keeps each range separate and ensures that this contamination does not occur.
+The field-theory Hamiltonian then takes the form
+$\mathcal{H}(x)=\begin{bmatrix}iu&D_{x}^{\dagger}\\ D_{x}&-iu\end{bmatrix},\qquad D_{x}=v(\partial_{x}+1)^{\alpha-1}-w(\partial_{x}+1)^{\alpha}.$ (S66)
+As on the lattice, the eigenvalue equations decouple at $E=\pm iu$. For a left-edge mode we set $E=iu$ and $\psi_{B}\equiv 0$, which gives the scalar boundary equation,
+$(\partial_{x}+1)^{\alpha-1}\,(v-w-w\,\partial_{x})\,\psi_{A}(x)=0.$ (S67)
+We now substitute the exponential ansatz $\psi_{A}(x)\sim e^{\beta x}$ into Eq. (S67), noting that only solutions with $\beta<0$ are normalizable.
+$(\beta+1)^{\alpha-1}\,\bigl[(v-w)-w\,\beta\bigr]=0,$ (S68)
+whose roots are
+$\beta=-1\quad(\text{$\alpha-1$-fold}),\qquad\beta=\frac{v-w}{w}.$ (S69)
+It follows immediately that the number of edge mode pairs is $\alpha-1$ for $v>w$ and $\alpha$ for $w>v$, in agreement with the lattice result.
+The $(\alpha-1)$-fold solution connects directly to the lattice picture.
+In the $\alpha$-nH SSH chain, the minimal hopping leg spans $\alpha$ sites; with an open boundary this leaves $\alpha-1$ sites that decouple from the bulk and thus form edge modes.
+In the continuum description, the lattice translation factor $e^{-i(\alpha-1)k}$ maps to $(\partial_{x}+1)^{\alpha-1}$, yielding edge solutions that mirror the same physics.
+After removing this factor, the residual equation
+$\bigl[(v-w)-w\,\partial_{x}\bigr]\psi_{A}=0$ (S70)
+is precisely the generalized mass-inversion condition discussed in the main text.
+To examine whether a kinetic-inversion–type picture applies, we consider the $\alpha=2$ case. The edge equation is
+$\bigl[-w\,\partial_{x}^{2}+(v-2w)\,\partial_{x}+(v-w)\bigr]\psi_{A}=0.$ (S71)
+On the line $v=w$, this reduces to the kinetic-inversion case discussed previously [78].
+Away from that line, however, the existence of an edge mode is not captured by a single sign flip (mass or kinetic inversion): for example, when $v>2w$ a negative root $\beta=-1$ still persists even though neither the effective mass nor the kinetic coefficient changes sign at the boundary.
+Because the second-order ordinary differential equation involves multiple independent coefficients, a one-parameter inversion picture is too restrictive; the appropriate criterion is the full characteristic equation.
+VI Generalized mass inversion and edge-state pinning at the interface
+Recall the field-theory equations in the nH bulk
+$(-\partial_{x}+m_{2})\psi_{B}(x)$ $=(E-iu)\psi_{A}(x),$
+$(\partial_{x}+m_{2})\psi_{A}(x)$ $=(E+iu)\psi_{B}(x),$ (S72)
+and in the Hermitian bulk
+$(-\partial_{x}+m_{1})\psi_{B}(x)$ $=E\psi_{A}(x),$
+$(\partial_{x}+m_{1})\psi_{A}(x)$ $=E\psi_{B}(x).$ (S73)
+We work at the topological QCP on the non-Hermitian side, $u=m_{2}$.
+Assume exponential behavior $\psi_{A,B}\sim e^{\kappa_{2}x}$ for $x>0$ (non-Hermitian side) and $\psi_{A,B}\sim e^{-\kappa_{1}x}$ for $x<0$ (Hermitian side); an interface-pinned mode then requires $\kappa_{1,2}<0$.
+This leads to the sublattice ratios
+$\frac{\psi_{A}}{\psi_{B}}$ $=\frac{E+iu}{\kappa_{2}+m_{2}}=\frac{-\kappa_{2}+m_{2}}{E-iu},$
+$\frac{\psi_{A}}{\psi_{B}}$ $=\frac{E}{-\kappa_{1}+m_{1}}=\frac{\kappa_{1}+m_{1}}{E}.$ (S74)
+On both sides the bulk dispersion relations are
+$m_{1}^{2}-\kappa_{1}^{2}$ $=E^{2},$
+$m_{2}^{2}-\kappa_{2}^{2}$ $=E^{2}+u^{2}.$ (S75)
+With $u=m_{2}$ this gives
+$\kappa_{2}$ $=iE\quad(\text{choosing the branch consistent with decay}),$
+$\kappa_{1}$ $=-\sqrt{m_{1}^{2}-E^{2}}.$ (S76)
+Continuity of $\psi_{A}/\psi_{B}$ at the interface requires
+$E(-\kappa_{2}+m_{2})=(E-iu)(\kappa_{1}+m_{1}).$ (S77)
+Substituting (S76) yields
+$E(-iE+m_{2})$ $=(E-iu)\bigl(-\sqrt{m_{1}^{2}-E^{2}}+m_{1}\bigr),$
+$E(-iE+m_{2})-(E-iu)m_{1}$ $=-(E-iu)\sqrt{m_{1}^{2}-E^{2}}$ (S78)
+Squaring both sides and simplifying gives
+$i4E^{3}u+i2E(E^{2}+u^{2})m_{1}=0,$ (S79)
+which leads to
+$E=\sqrt{\frac{u^{2}m_{1}}{2u-m_{1}}}=iu\sqrt{\frac{-m_{1}}{2u-m_{1}}}.$ (S80)
+From the solution we can write $E=iau$. For $m_{1}<0$ one finds $a\in[0,1]$, giving an interface edge mode (mass inversion). As $m_{1}\to 0$, the mode merges into the bulk on both sides.
+By contrast, when $m_{1}>2u$ we have
+$a=\sqrt{\frac{m_{1}}{m_{1}-2u}}>1.$ (S81)
+Substituting $E=iau$ and the $\kappa$ solutions (S76) into the sublattice ratios (S74) gives
+$\left(\frac{\psi_{A}}{\psi_{B}}\right)_{R}$ $=\frac{iau+iu}{-au+u}=i\frac{a+1}{1-a}=-i\frac{a+1}{a-1}\quad(<0\ \text{imag}),$
+$\left(\frac{\psi_{A}}{\psi_{B}}\right)_{L}$ $=\frac{iau}{\sqrt{m_{1}^{2}+a^{2}u^{2}}+m_{1}}=i\frac{au}{\sqrt{m_{1}^{2}+a^{2}u^{2}}+m_{1}}\quad(>0\ \text{imag}).$ (S82)
+Thus the right ratio is purely negative imaginary while the left ratio is purely positive imaginary; they cannot match, so the boundary condition fails.
+This shows that the $m_{1}>2u$ branch (with $a>1$) is an extraneous root generated by squaring the equation.
+Interface edge modes occur only when mass inverse at the interface.
+These analytic results are borne out by the numerics.
+In the limit $m_{1}\to$-$\infty$, Eq.
+(S80) yields $E=iu$.
+From (S76) and (S74), the interface mode has decay rates $\kappa_{2}=-u$ and $\kappa_{1}\to$-$\infty$, with a vanishing $B$-sublattice amplitude.
+The solution therefore reduces to the OBC edge pinned on the non-Hermitian side, in agreement with the numerical profile shown in Fig.
+S8(a).
+Physically, this is natural—the infinite-mass boundary is equivalent to vacuum.
+For finite negative $m_{1}$, the edge mode remains pinned at the interface but now with a finite decay length on both sides, which is visible in Fig.
+S8(b).
+An additional feature appears here: the plotted $B$-sublattice amplitude is negative.
+This can be explained by the imaginary sublattice ratio $\psi_{A}/\psi_{B}\sim+i$ as discussed above.
+The amplitude is computed in the biorthogonal basis ($\psi_{L}^{*}\psi_{R}$, effectively $\psi^{2}$ here), thus we obtain a negative $B$-sublattice amplitude.
+The sign structure is physically consistent: the on-site imaginary potentials have opposite signs on $A/B$, so for a positive imaginary energy the negative $B$-sublattice amplitude aligns with the negative imaginary $B$-sublattice potential in the biorthogonal expectation, yielding a net positive imaginary contribution.
+If the configuration is reversed (non-Hermitian part on the left, Hermitian on the right), the roles swap and the energy becomes negative imaginary, producing the opposite sign pattern.
+We can also recover the familiar representation by computing the unit cell amplitude (Fig.
+S8(c)), which reproduces the plot shown in the main text Fig.
+4(b).
+Finally, as $m_{1}\to 0^{-}$ the analytic solution enforces $\kappa_{1,2}\to 0^{-}$ and $E\to 0$, so the edge state merges into the bulk on both sides.
+The numerical profiles in Fig.
+S8(d) confirm this: the decay lengths diverge on both sides.
+This contrasts with a purely Hermitian interface, where the decay length is set solely by local parameters; here, via $E$ and (S76), the two bulks are coupled, so localization depends on both sides of the interface.
+In summary, although the simple decoupling argument no longer applies at an interface, the generalized mass-inversion picture remains valid and continues to govern the existence of pinned edge states.
+Refer to caption(a)
+Refer to caption(b)
+Refer to caption(c)
+Refer to caption(d)
+Refer to caption(e)
+[…]
+Lattice calculation—For completeness, the same problem can be formulated directly in the lattice model.
+We set up the lattice interface as in Fig.
+S8(e): the left side is the Hermitian SSH chain, and the right side is the nH SSH chain at its topological QCP.
+We label the unit cell adjacent to the interface by $i$.
+The single-particle equations near the interface are
+$-w\psi_{i-1}^{B}+v_{1}\psi_{i}^{B}$ $=E\psi_{i}^{A},$
+$v_{1}\psi_{i}^{A}-w\psi_{i+1}^{A}$ $=E\psi_{i}^{B},$
+$-w\psi_{i}^{B}+v_{2}\psi_{i+1}^{B}$ $=(E-iu)\psi_{i+1}^{A},$
+$v_{2}\psi_{i+1}^{A}-w\psi_{i+2}^{A}$ $=(E+iu)\psi_{i+1}^{B},$ (S83)
+In each bulk we adopt an exponential ansatz with a fixed sublattice ratio:
+$\frac{\psi^{B}_{n}}{\psi^{A}_{n}}$ $=\left(\frac{\psi^{B}}{\psi^{A}}\right)_{L}=\text{const.},$ $\qquad\psi^{A,B}_{n-1}$ $=\beta_{L}\psi^{A,B}_{n},$ $\qquad n$ $\in\text{left},$
+$\frac{\psi^{B}_{n}}{\psi^{A}_{n}}$ $=\left(\frac{\psi^{B}}{\psi^{A}}\right)_{R}=\text{const.},$ $\qquad\psi^{A,B}_{n+1}$ $=\beta_{R}\psi^{A,B}_{n},$ $\qquad n$ $\in\text{right}.$ (S84)
+For an edge mode pinned at the interface we require $|\beta_{1,2}|<1$. Substitute the ansatz into the bulk eigen–equations to obtain the recursion relations
+$\left(\frac{\psi^{B}}{\psi^{A}}\right)_{L}$ $=\frac{E}{v_{1}-w\beta_{L}}=\frac{v_{1}-w\beta_{L}^{-1}}{E},$
+$\left(\frac{\psi^{B}}{\psi^{A}}\right)_{R}$ $=\frac{E-iu}{v_{2}-w\beta_{R}^{-1}}=\frac{v_{2}-w\beta_{R}}{E+iu}.$ (S85)
+Equating the two forms on each side, these can be rearranged into the bulk dispersions
+$E^{2}$ $=v_{1}^{2}+w^{2}-v_{1}w\left(\beta_{L}+\beta_{L}^{-1}\right),$
+$E^{2}+u^{2}$ $=v_{2}^{2}+w^{2}-v_{2}w\left(\beta_{R}+\beta_{R}^{-1}\right).$ (S86)
+These relations are the lattice counterparts of the continuum bulk dispersions (S76).
+For the interface boundary condition, we start from the eigen-equations (S83) across the interface and substitute the fixed sublattice ratios to obtain
+$-w\psi_{i+1}^{A}$ $=\Big[E-v_{1}\left(\tfrac{\psi^{A}}{\psi^{B}}\right)_{L}\Big]\psi_{i}^{B},$
+$-w\psi_{i}^{B}$ $=\Big[E-iu-v_{2}\left(\tfrac{\psi^{B}}{\psi^{A}}\right)_{R}\Big]\psi_{i+1}^{A}.$ (S87)
+Eliminating the ratio gives the boundary-matching condition:
+$\frac{E-v_{1}\left(\tfrac{\psi^{A}}{\psi^{B}}\right)_{L}}{-w}=\frac{-w}{E-iu-v_{2}\left(\tfrac{\psi^{B}}{\psi^{A}}\right)_{R}}.$ (S88)
+In the long-wavelength limit, the sublattice ratio varies negligibly across a unit cell, so the lattice condition (S88) reduces to the continuum boundary condition used in the field-theory treatment.
+Upon here, we have all the equations needed to solve the interface problem: together, Eqs.
+(S86) and (S88) provide three independent equations for the three unknowns $E,\beta_{L},\beta_{R}$.
+However, while the ansatz is simple, solving explicitly for $\beta_{L,R}$ and $E$ leads to cumbersome algebra with no clean closed form.
+We therefore present the lattice formulation solely as a reference and numerical validation of the continuum analysis.